# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

### v2

@@ 1 Introduction
Large Language Model (LLM) agents have rapidly transitioned from proof-of-concept chatbots to end-to-end systems that can write software (Wang et al., 2024c), control browsers (Müller & Žunič, 2024), and reason over multi-modal inputs. Frameworks such as Manus, OWL (Hu et al., 2025), OpenHands (Wang et al., 2024c), and Codex routinely solve complex, tool-rich tasks and achieve state-of-the-art results on agentic benchmarks like GAIA (Mialon et al., 2023) and SWE-Bench (Jimenez et al., 2023). Yet these evaluations focus almost exclusively on reasoning (planning, tool using, code synthesis) and leave the equally important question of memorization (abstraction, storing, updating, retrieving) largely under-explored. Recent memory-centric architectures—ranging from parametric memory systems like MemoryLLM (Wang et al., 2024d), SELF-PARAM (Wang et al., ), and M+(Wang et al., 2025) to commercial token-level memory solutions such as MemGPT(Packer et al., 2023; Lin et al., 2025), Mem0(Chhikara et al., 2025), Cognee(Markovic et al., 2025), and Zep(Rasmussen et al., 2025)—employ diverse strategies for storing and retrieving past information. Despite growing interest, their real-world effectiveness remains largely anecdotal, and there is currently no unified benchmark for systematically evaluating the quality of memory in agents. In this paper, we refer to agents equipped with memory mechanisms as Memory Agents, where memory can take various forms, including parameters, vectors, textual histories, or external databases. In this paper, we primarily focus on memory agents that utilize textual histories and external databases, as these approaches are most commonly deployed in real-world applications. In contrast, memory encoded in model parameters (Wang et al., 2024d; 2025; Yin et al., 2024) remains largely within academic research and is typically less capable than proprietary memory systems equipped on closed-sourced API models.
@@
Based on some classic theories in memory and cognitive science (James, 1890; McClelland et al., 1995; Anderson & Neely, 1996; Wimber et al., 2015), we identify four complementary competencies (Examples shown in Figure 1) to evaluate memory agents: (1) Accurate Retrieval (AR): The ability to extract the correct snippet in response to a query. This can involve one-hop or multi-hop retrieval, as long as the relevant information can be accessed with a single query. (2) Test-Time Learning (TTL): The capacity to incorporate new behaviors or acquire new skills during deployment, without additional training. (3) Long-Range Understanding (LRU): The ability to integrate information distributed across extended contexts ( 100k tokens) and answer questions requiring a global understanding of the entire sequence. (4) Selective Forgetting (SF): The skill to revise, overwrite, or remove previously stored information when faced with contradictory evidence, aligning with goals in model editing and knowledge unlearning tasks (Meng et al., 2023; Wang et al., 2024e). For these fuor competencies, we provide more detailed definitions in Appendix A.
@@
To address these limitations, we introduce a unified benchmark framework, MemoryAgentBench, specifically designed to evaluate a broad spectrum of memory mechanisms in agent systems. We also provide a framework for memory agent evaluation. In this framework, agents are presented with sequences of textual inputs that simulate multi-turn interactions with users. We reconstructed existing datasets originally developed for long-context LLM evaluation by segmenting and reconstructing inputs into multiple dialogue chunks and feeding them incrementally to the agent in a time order. However, since these datasets do not fully capture all four targeted memory competencies, we also introduce two new datasets: EventQA and FactConsolidation, designed to evaluate accurate retrieval and selective forgetting, respectively. Our benchmark includes evaluations of state-of-the-art commercial memory agents (such as MIRIX and MemGPT), long-context agents that treat the full input as memory, and RAG agents that extend their memory through retrieval methods. We examine how techniques developed for long-context models and RAG transfer to the memory agent setting. By providing a consistent evaluation protocol across diverse agent architectures and datasets, MemoryAgentBench delivers comprehensive insights into agent performance across the four core memory competencies.
@@
Empirical Study: We implement various simple agents with diverse memory mechanisms, adopt commercial agents, and evaluate these agents on our proposed benchmark. With our results, we show that existing memory agents, while effective in some tasks, still face significant challenges on some aspects.
@@ 2 Related Work
2.1 Benchmarks with Long Input
In this section, we review prior work on long-context benchmarks. Early benchmarks designed for long-context evaluation include LongBench(Bai et al., 2023) and LooGLE(Li et al., 2023), with average input lengths of approximately 20k and 24k tokens, respectively. More recent benchmarks—such as -Bench (Zhang et al., 2024), HELMET(Yen et al., 2024), RULER(Hsieh et al., 2024), NOCHA(Karpinska et al., 2024), NoLiMa (Modarressi et al., 2025) and LongBench V2(Bai et al., 2024)—extend context lengths to over 100k tokens and are primarily intended to evaluate the capabilities of long-context models. However, despite their scale, these benchmarks are not designed to assess memory agents, and no prior work has repurposed them for that goal. More recently, LOCOMO (Maharana et al., 2024), LongMemEval (Wu et al., 2025), RealTalk (Lee et al., 2025) and StoryBench (Wan & Ma, 2025) have been proposed specifically for evaluating memory agents. While promising, LOCOMO still features relatively short conversations (9k), and LongMemEval uses synthetic conversations with limited topical diversity, making the dialogues less realistic and potentially less representative of real-world memory use cases. Meanwhile, the evaluation scope of the above benchmarks is not sufficient to comprehensively assess long-term memory from multiple dimensions.
@@ 2.2 Agents with Memory Mechanisms
Memory mechanisms are attracting more and more attention lately (Wang et al., 2025/02). Recent advancements in LLMs have demonstrated the capability to process extended context lengths, ranging from 100K to over 1 million tokens. For instance, models such as GPT-4o (OpenAI, 2025b) and Claude 3.7 (Anthropic, 2025) can handle inputs of approximately 100K to 200K tokens, while models like Gemini 2.0 Pro (DeepMind, 2025) and the GPT-4.1 series extend this capacity beyond 1 million tokens. These strong long-context capabilities enable a simple yet effective form of memory: storing information directly within the context window. However, this approach is inherently constrained by a hard limit—once the context window is exceeded, earlier information must be discarded.
@@
In parallel, RAG continues to serve as a dominant paradigm for managing excessive context. By retrieving relevant information from earlier context and feeding it to the LLM, RAG allows systems to overcome context length limitations. For example, OpenAI’s recent memory functionality111https://openai.com/index/memory-and-new-controls-for-chatgpt/ combines explicit user preference tracking with retrieval-based methods that reference prior interactions. RAG methods can be broadly classified into three categories: Simple RAG: These methods rely on string-matching techniques such as TF-IDF, BM25 (Robertson & Walker, 1994), and BMX (Li et al., 2024), which are entirely non-neural and operate on string-level similarity. Embedding-based RAG: This class leverages neural encoders, primarily transformers, to map text into dense vector representations (Wu et al., 2022). Early methods like DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2021) are based on BERT (Devlin et al., 2019), while more recent models such as Qwen3-Embedding (Zhang et al., 2025) achieve significantly improved retrieval performance. Structure-Augmented RAG: These approaches enhance retrieval with structural representations such as graphs or trees. Representative systems include GraphRAG (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), HippoRAG-V2 (Gutiérrez et al., 2025), Cognee, Zep (Rasmussen et al., 2025), MemoRAG (Qian et al., 2025), Mem0 (Chhikara et al., 2025), MemoryOS (Kang et al., 2025), Memary (kingjulio8238 & Memary contributors, 2024) and Memobase (memodb-io & Memobase contributors, 2025). Despite their effectiveness, RAG-based methods face challenges with ambiguous queries, multi-hop reasoning, and long-range comprehension. When questions require integrating knowledge across an entire session or learning from long, skill-encoding inputs, the retrieval mechanism—limited to the top-k most relevant passages—may fail to surface the necessary information. To address these limitations, Agentic Memory Agents introduce an iterative, decision-driven framework. Rather than relying on a single-pass retrieval, these agents dynamically process the query, retrieve evidence, reflect, and iterate through multiple retrieval and reasoning cycles. Examples include MemGPT (Packer et al., 2023), Self-RAG (Asai et al., 2023), Auto-RAG (Yu et al., 2024), A-MEM (Xu et al., 2025), Mem1 (Zhou et al., 2025), MemAgent (Yu et al., 2025), and MIRIX (Wang & Chen, 2025). This agentic design is particularly effective for resolving ambiguous or multi-step queries. Nonetheless, these methods remain fundamentally constrained by the limitations of RAG—namely, the inability to fully understand or learn from long-range context that is inaccessible via retrieval alone.
@@ 3.1 Dataset Preperation
In this section, we describe how we reconstruct existing datasets and build new ones for evaluating each competency aspect. All datasets with their categories are shown in Table 1. We introduce the details in datasets curation in Appendix A.
@@
We evaluate TTL via two task categories: (1) Multi-Class Classification (MCC): We reconstructed five classification datasets used in prior TTL work (Bertsch et al., 2024; Yen et al., 2024): BANKING77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), TREC-Coarse, TREC-Fine (Li & Roth, 2002), and NLU (Liu et al., 2019). Each task requires the agent to map sentences to class labels, leveraging previously seen labeled examples in context. (2) Recommendation: Based on the setup from (Li et al., 2018; He et al., 2023), we construct a dataset to evaluate movie recommendation via dialogue history. The agent is exposed to thousands of movie-related dialogue turns and is asked to recommend twenty relevant movies based on the long interaction history.
@@
To assess whether an agent can forget out of date memory and reason over them, we construct a new dataset called FactConsolidation. Specifically, We build this benchmark using counterfactual edit pairs from MQUAKE (Zhong et al., 2023). Each pair contains a true fact and a rewritten, contradictory version. These are ordered such that the rewritten (new) fact appears after the original, simulating a realistic update scenario. We concatenate multiple such edit pairs to create long contexts of length 6K, 32K, 64K, 262K. We then adpot MQUAKE’s original questions and categorize them into: (1) FactConsolidation-SH (Ours) (SH means Single-Hop), requiring direct factual recall (e.g., “Which country was tool created in?”), and (2) FactConsolidation-MH (Ours) (MH refers to Multi-Hop), requiring inference over multiple facts (e.g., “What is the location of death of the spouse of person ?”). Agents are prompted to prioritize later information in case of conflict and reason based on the final memory state. This setup directly evaluates the strength and consistency of selective forgetting over long sequences.
@@ 3.3 Datasets and Agents Formulation
We standardize all datasets into the format: (chunks), (questions), and (answers), where denotes the -th chunk wrapped to construct a user message with instructions of memorizing the content in a sequential input, and represents a single conversation. Each chunk is accompanied by instructions prompting the agent to memorize its contents. Example prompts are provided in Appendix C.1. When curating datasets like EventQA and FactConsolidation, we deliberately design scenarios where multiple questions follow a single context. This allows us to probe the model’s memory multiple times with one sequential injection. For example, in LME (S*), five contexts are paired with 300 questions (shown in Table 5 in Appendix A). This design choice reflects a key trend: as LLMs support increasingly long context windows and memory agents become more capable of handling extended inputs, evaluation datasets must also scale accordingly. Injecting 1M tokens for just one question is resource-inefficient, whereas associating the same input with many questions provides significantly higher utility.
@@
In our framework, all agents are required to take the chunks one by one, absorb them into memory, and incrementally update the memory. After seeing all the chunks, we ask the agent to answer the related questions.

### v3

@@ 1 Introduction
Large Language Model (LLM) agents have rapidly transitioned from proof-of-concept chatbots to end-to-end systems that can write software (Wang et al., 2024c), control browsers (Müller and Žunič, 2024), and reason over multi-modal inputs. Frameworks such as Manus, OWL (Hu et al., 2025), OpenHands (Wang et al., 2024c), and Codex routinely solve complex, tool-rich tasks and achieve state-of-the-art results on agentic benchmarks like GAIA (Mialon et al., 2023) and SWE-Bench (Jimenez et al., 2023). Yet these evaluations focus almost exclusively on reasoning (planning, tool using, code synthesis) and leave the equally important question of memorization (abstraction, storing, updating, retrieving) largely under-explored. Recent memory-centric architectures—ranging from parametric memory systems like MemoryLLM (Wang et al., 2024d), SELF-PARAM (Wang et al., ), and M+ (Wang et al., 2025b) to commercial token-level memory solutions such as MemGPT (Packer et al., 2023; Lin et al., 2025), Mem0 (Chhikara et al., 2025), Cognee (Markovic et al., 2025), Zep (Rasmussen et al., 2025) and MIRIX (Wang and Chen, 2025)—employ diverse strategies for storing and retrieving past information. Despite growing interest, their real-world effectiveness remains largely anecdotal, and there is currently no unified benchmark for systematically evaluating the quality of memory in agents. In this paper, we refer to agents equipped with memory mechanisms as Memory Agents, where memory can take various forms, including parameters, vectors, textual histories, or external databases. In this paper, we primarily focus on memory agents that utilize textual histories and external databases, as these approaches are most commonly deployed in real-world applications. In contrast, memory encoded in model parameters (Wang et al., 2024d; 2025b; Yin et al., 2024) remains largely within academic research and is typically less capable than proprietary memory systems equipped on closed-sourced API models.
@@
Based on some classic theories in memory and cognitive science (James, 1890; McClelland et al., 1995; Anderson and Neely, 1996; Wimber et al., 2015), we identify four complementary competencies (Examples shown in Figure 1) to evaluate memory agents: (1) Accurate Retrieval (AR): The ability to extract the correct snippet in response to a query. This can involve one-hop or multi-hop retrieval, as long as the relevant information can be accessed with a single query. (2) Test-Time Learning (TTL): The capacity to incorporate new behaviors or acquire new skills during deployment, without additional training. (3) Long-Range Understanding (LRU): The ability to integrate information distributed across extended contexts ( 100k tokens) and answer questions requiring a global understanding of the entire sequence. (4) Selective Forgetting (SF): The skill to revise, overwrite, or remove previously stored information when faced with contradictory evidence, aligning with goals in model editing and knowledge unlearning tasks (Meng et al., 2023; Wang et al., 2024e). For these four competencies, we provide more detailed definitions in Appendix B.
@@
To address these limitations, we introduce a unified benchmark, MemoryAgentBench, specifically designed to evaluate a broad spectrum of memory mechanisms in agent systems. We also provide a framework for memory agent evaluation. In this framework, agents are presented with sequences of textual inputs that simulate multi-turn interactions with users. We reconstructed existing datasets originally developed for long-context LLM evaluation by segmenting and reconstructing inputs into multiple dialogue chunks and feeding them incrementally to the agent in a time order. However, since these datasets do not fully capture all four targeted memory competencies, we also introduce two new datasets: EventQA and FactConsolidation, designed to evaluate accurate retrieval and selective forgetting, respectively. Our benchmark includes evaluations of state-of-the-art commercial memory agents (such as MIRIX and MemGPT), long-context agents that treat the full input as memory, and RAG agents that extend their memory through retrieval methods. We examine how techniques developed for long-context models and RAG transfer to the memory agent setting. By providing a consistent evaluation protocol across diverse agent architectures and datasets, MemoryAgentBench delivers comprehensive insights into agent performance across the four core memory competencies. In Table 1, we compare the MemoryAgentBench with previous representative benchmarks across multiple dimensions.
@@
Empirical Study: We implement various simple agents with diverse memory mechanisms, adopt commercial agents, and evaluate these agents on our proposed benchmark. Our results demonstrate that existing memory agents, although effective in some tasks, still face significant challenges in certain aspects.
@@ 2 Related Work
2.1 Benchmarks on Long-Context and Memory
In this section, we review prior work on evaluation benchmarks, categorizing them into three domains: long-context understanding, retrieval-augmented generation, and memory agents.
Benchmarks for Long-Context LLMs.
Early benchmarks designed for long-context evaluation include LongBench(Bai et al., 2023) and LooGLE(Li et al., 2023), with average input lengths of approximately 20k and 24k tokens, respectively. More recent benchmarks—such as -Bench (Zhang et al., 2024), HELMET(Yen et al., 2024), RULER(Hsieh et al., 2024), NOCHA(Karpinska et al., 2024), NoLiMa (Modarressi et al., 2025) and LongBench V2(Bai et al., 2024)—extend context lengths to over 100k tokens. While these benchmarks effectively assess the model’s ability to process extensive information in a single pass, they are primarily intended for static long-context reading comprehension and do not reflect the incremental, multi-turn nature of memory agents.
Benchmarks for Retrieval-Augmented Generation.
Beyond pure long-context evaluation, a line of benchmarks targets retrieval-augmented generation (RAG) for knowledge-intensive tasks such as open-domain QA, fact checking, and document ranking over fixed corpora, e.g., KILT (Petroni et al., 2021) and BEIR (Thakur et al., 2021). More recent work explicitly evaluates end-to-end RAG systems under long-context or application-specific scenarios, including LaRA (Li et al., ), LONG2RAG (Qi et al., 2024), FRAMES (Krishna et al., 2025), and CRUD-RAG (Lyu et al., 2025). Large-scale benchmarks such as RAGBench (Friel et al., 2024), RAGTruth (Niu et al., 2024), FreshLLM (Vu et al., 2024), and -RAGBench (Strich et al., 2025) further extend the evaluation space to industrial manuals, hallucination detection, time-sensitive web QA, and text-and-table financial reports, respectively. However, existing RAG benchmarks typically assume a static or slowly changing knowledge base and short-lived interactions, emphasizing retrieval accuracy and grounding over the continual updating and selective forgetting of information that is central to memory agents.
Benchmarks for Memory Agents
More recently, benchmarks such as LOCOMO (Maharana et al., 2024), LongMemEval (Wu et al., 2025), RealTalk (Lee et al., 2025) and StoryBench (Wan and Ma, 2025) have been proposed specifically for evaluating memory agents. While promising, LOCOMO still features relatively short conversations (9k), and LongMemEval uses synthetic conversations with limited topical diversity, making the dialogues less realistic and potentially less representative of real-world memory use cases. Meanwhile, the evaluation scope of the above benchmarks is not sufficient to comprehensively assess the four core competencies—Accurate Retrieval, Test-Time Learning, Long-Range Understanding, and Selective Forgetting—that are essential for robust memory agents.
| Benchmark | #Q | Context Depth | Core Memory Competencies | Evaluation Coverage | |||||
| AR | TTL | LRU | SF | LCA | RAG | AM | |||
| MemoryBank (Zhong et al., 2023a) | 194 | 5k | ✓ | ✓ | ✗ | ✗ | ✓ | ✗ | ✓ |
| LoCoMo (Maharana et al., 2024) | 7512 | 10k | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✗ |
| PerLTQA (Du et al., 2024) | 8593 | 1M* | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✗ |
| RealTalk (Lee et al., 2025) | 728 | 375k* | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| LongMemEval (Wu et al., 2025) | 500 | 115k, 1.5M | ✓ | ✗ | ✗ | ✗ | ✓ | ✓ | ✗ |
| StoryBench (Wan and Ma, 2025) | 86 | - | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✗ |
| MemoryAgentBench | 2071 | 103k-1.44M | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@ 2.2 Agents with Memory Mechanisms
Memory mechanisms are attracting more and more attention lately (Wang et al., 2025a). Recent advancements in LLMs have demonstrated the capability to process extended context lengths, ranging from 100K to over 1 million tokens. For instance, models such as GPT-4o (OpenAI, 2025a) and Claude 3.7 (Anthropic, 2025) can handle inputs of approximately 100K to 200K tokens, while models like Gemini 2.0 Pro (DeepMind, 2025) and the GPT-4.1 series extend this capacity beyond 1 million tokens. These strong long-context capabilities enable a simple yet effective form of memory: storing information directly within the context window. However, this approach is inherently constrained by a hard limit—once the context window is exceeded, earlier information must be discarded.
@@
In parallel, RAG continues to serve as a dominant paradigm for managing excessive context. By retrieving relevant information from earlier context and feeding it to the LLM, RAG allows systems to overcome context length limitations. For example, OpenAI’s recent memory functionality111https://openai.com/index/memory-and-new-controls-for-chatgpt/ combines explicit user preference tracking with retrieval-based methods that reference prior interactions. RAG methods can be broadly classified into three categories: Simple RAG: These methods rely on string-matching techniques such as TF-IDF, BM25 (Robertson and Walker, 1994), and BMX (Li et al., 2024), which are entirely non-neural and operate on string-level similarity. Embedding-based RAG: This class leverages neural encoders, primarily transformers, to map text into dense vector representations (Wu et al., 2022). Early methods like DPR (Karpukhin et al., 2020) and Contriever (Izacard et al., 2021) are based on BERT (Devlin et al., 2019), while more recent models such as Qwen3-Embedding (Zhang et al., 2025) achieve significantly improved retrieval performance. Structure-Augmented RAG: These approaches enhance retrieval with structural representations such as graphs or trees. Representative systems include GraphRAG (Edge et al., 2024), RAPTOR (Sarthi et al., 2024), HippoRAG-V2 (Gutiérrez et al., 2025), Cognee, Zep (Rasmussen et al., 2025), MemoRAG (Qian et al., 2025), Mem0 (Chhikara et al., 2025), MemoryOS (Kang et al., 2025), Memary (kingjulio8238 and Memary contributors, 2024) and Memobase (memodb-io and Memobase contributors, 2025). Despite their effectiveness, RAG-based methods face challenges with ambiguous queries, multi-hop reasoning, and long-range comprehension. When questions require integrating knowledge across an entire session or learning from long, skill-encoding inputs, the retrieval mechanism—limited to the top-k most relevant passages—may fail to surface the necessary information. To address these limitations, Agentic Memory Agents introduce an iterative, decision-driven framework. Rather than relying on a single-pass retrieval, these agents dynamically process the query, retrieve evidence, reflect, and iterate through multiple retrieval and reasoning cycles. Examples include MemGPT (Packer et al., 2023), Self-RAG (Asai et al., 2023), Auto-RAG (Yu et al., 2024), A-MEM (Xu et al., 2025), Mem1 (Zhou et al., 2025), MemAgent (Yu et al., 2025), and MIRIX (Wang and Chen, 2025). This agentic design is particularly effective for resolving ambiguous or multi-step queries. Nonetheless, these methods remain fundamentally constrained by the limitations of RAG—namely, the inability to fully understand or learn from long-range context that is inaccessible via retrieval alone.
@@ 3.1 Dataset Preperation
In this section, we describe how we reconstruct existing datasets and build new ones for evaluating each competency aspect. All datasets with their categories are shown in Table 2. We introduce the details in datasets curation in Appendix B.
@@
We evaluate TTL via two task categories: (1) Multi-Class Classification (MCC): We reconstructed five classification datasets used in prior TTL work (Bertsch et al., 2024; Yen et al., 2024): BANKING77 (Casanueva et al., 2020), CLINC150 (Larson et al., 2019), TREC-Coarse, TREC-Fine (Li and Roth, 2002), and NLU (Liu et al., 2019). Each task requires the agent to map sentences to class labels, leveraging previously seen labeled examples in context. (2) Recommendation: Based on the setup from (Li et al., 2018; He et al., 2023), we construct a dataset to evaluate movie recommendation via dialogue history. The agent is exposed to thousands of movie-related dialogue turns and is asked to recommend twenty relevant movies based on the long interaction history.
@@
To assess whether an agent can forget out of date memory and reason over them, we construct a new dataset called FactConsolidation. Specifically, We build this benchmark using counterfactual edit pairs from MQUAKE (Zhong et al., 2023b). Each pair contains a true fact and a rewritten, contradictory version. These are ordered such that the rewritten (new) fact appears after the original, simulating a realistic update scenario. We concatenate multiple such edit pairs to create long contexts of length 6K, 32K, 64K, 262K. We then adpot MQUAKE’s original questions and categorize them into: (1) FactConsolidation-SH (Ours) (SH means Single-Hop), requiring direct factual recall (e.g., “Which country was tool created in?”), and (2) FactConsolidation-MH (Ours) (MH refers to Multi-Hop), requiring inference over multiple facts (e.g., “What is the location of death of the spouse of person ?”). Agents are prompted to prioritize later information in case of conflict and reason based on the final memory state. This setup directly evaluates the strength and consistency of selective forgetting over long sequences.
@@ 3.3 Datasets and Agents Formulation
We standardize all datasets into the format: (chunks), (questions), and (answers), where denotes the -th chunk wrapped to construct a user message with instructions of memorizing the content in a sequential input, and represents a single conversation. Each chunk is accompanied by instructions prompting the agent to memorize its contents. Example prompts are provided in Appendix D.1. When curating datasets like EventQA and FactConsolidation, we deliberately design scenarios where multiple questions follow a single context. This allows us to probe the model’s memory multiple times with one sequential injection. For example, in LME (S*), five contexts are paired with 300 questions (shown in Table 6 in Appendix B). This design choice reflects a key trend: as LLMs support increasingly long context windows and memory agents become more capable of handling extended inputs, evaluation datasets must also scale accordingly. Injecting 1M tokens for just one question is resource-inefficient, whereas associating the same input with many questions provides significantly higher utility.
Prompt Formulation and Interaction Protocol
Unlike standard long-context evaluations that input raw text, we wrap all input chunks within a simulated User-Assistant dialogue to explicitly trigger the agent’s memory mechanism. Each input chunk is preceded by a memorization instruction (e.g., “Please memorize it and I will ask some questions…”) to establish a clear intent for information storage. Simultaneously, for each specific dataset, we carefully curated the instructions to ensure agents accurately comprehend the task intent and execute the required actions. Crucially, for the Selective Forgetting competency, we introduce explicit guardrails in the prompt. We explicitly instruct agents that facts are indexed by serial numbers, and that “newer facts have larger serial numbers.”. The agents are mandated to solve conflicts by finding the newest fact (see Appendix D for full templates).
@@
In our framework, all agents are required to take the chunks one by one, absorb them into memory, and incrementally update the memory. After seeing all the chunks, we ask the agent to answer the related questions. To guarantee fair comparison, we employed standardized prompt templates across all agents within each evaluation category, with only minimal adaptations where necessary.